Two protons in an atomic nucleus are typically separated by a distance of 2 x 10-15 m. The electric repulsion force between the protons is huge, but the attractive nuclear force is even stronger and keeps the nucleus from bursting apart. What is the magnitude of the electric force between two protons separated by 2 x 10-15 m?


1
Expert's answer
2021-04-28T09:54:32-0400

The magnitude of the electric force between two protons can be found by Coulomb's law:


"F=\frac{e·e}{4\pi\epsilon_0 r^2}=57.7\text{ N}."